# Orthogonal projections of hypercubes

### v3

@@ VIII Conclusions
Our contributions are summarized as follows: (1) We introduce the Hamming and fractal projection of hypercubes in Sec. II. (2) We propose interpreting the biplot basis of PCA as the basis of the orthogonal projection of hypercubes in Sec. III. (3) We find that the PC1 loading of PCA on hypercubes is equivalent to the weighted hypercubic vertices in Sec. IV, and validate it analytically and numerically in Sec. V. (4) We reveal numerically and analytically that the quality of the orthogonal projection of hypercubes tends to be worse around the origin of the projected space in Sec. VI. (5) We visualize the hypercubic energy landscapes of finite artificial spin-ice systems and extract physical interpretations of them in Sec. VII.
Although we demonstrate that PCA produces reproducible, interpretable, and automatic projections of hypercubes, the remaining challenge is the overlapping of the hypercubic vertices in the projection. As we show numerically and analytically in Sec. VI, PCA locates numerous hypercubic vertices around the origin of the projected space, and those overlapped hypercubic vertices can increase the error of the projection. We, however, analytically know these limitations of PCA beforehand: we can interpret the resulting plot while grasping the limitations, unlike other nonlinear dimensionality reduction methods, where such analytical properties are not obvious and might be more difficult to understand. When we create the hypercubic energy landscape in Sec. VII, we avoid the overlapping problem by projecting a limited number of hypercubic vertices. With a similar strategy, one might bypass this difficulty in other applications.
H. S. M. Coxeter mentioned that illustrations of high-dimensional objects have psychological and artistic merit [10]. In this study, we initiate unveiling the new merit—or possibility—of projecting high-dimensional hypercubes: interpreting binary data through data-driven visualization. As Anscombe’s quartet [133] shows hidden patterns not in statistics but in graphs, informative orthogonal projection using PCA might lead to fresh understanding, interpretation, and discovery [134, 135] in high-dimensional binary data across sciences.
Calculations and visualizations of this work were performed using open-source python [136] libraries: matplotlib [137], networkx [138], numpy [139], and scipy [140]. All codes will be available online [141].
@@ Acknowledgements.
Special thanks go to Takashi Horiike, Kennichi Hagiwara, Nozomu Imai, Takumi Nishina, and Hiroto Murata for fruitful discussions at the early stage of this research. We thank valuable comments on this manuscript from Mizuho Aoki, Johanna von Drachenfels, Anton G. Golles, and Henri Schmidt. This work was supported by KAKENHI Grants 22H00406 and 24H00061 of the Japan Society for the Promotion of Science.
Y.H. performed conceptualization, formal analysis, software development, validation, investigation, data curation, and visualization; Y.H. and S.F. developed the methodology; S.F. provided computational resources; Y.H. wrote the original draft; and Y.H. and S.F. reviewed and edited the manuscript.
@@ Appendix A Constructing hypercubes
A hypercube is constructed as follows [10, 142, 13]. Suppose one has a zero-dimensional point, one obtains a one-dimensional line by moving the point in a direction. By moving the line in a direction but not along its line, one can obtain a two-dimensional parallelogram. Again, moving the parallelogram in a direction but not along its plane, one creates a three-dimensional parallelepiped. By repeating this translation process $N$ times, one obtains an $N$-dimensional parallelepiped or a parallelotope with $2^{N}$ vertices and $\frac{2^{N}N}{2}$ edges.111111 Each of the $2^{N}$ vertices has $N$ edges. To avoid double counting, one needs to divide them by two. The unit vectors of each dimension correspond to the direction of the translation, i.e., edges created by the translation are parallel to the directions of unit vectors.
@@ B.1 Maximum projection variance formulation
Multiplying $\bm{u}_{1}^{\top}$ to both sides of the equation from the left side yields121212 $\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{2}=\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{1}=\bm{u}_{2}^{\top}\lambda_{1}\bm{u}_{1}=\lambda_{1}\bm{u}_{2}^{\top}\bm{u}_{1}=0$ because of the orthogonality between $\bm{u}_{1}$ and $\bm{u}_{2}$, $\bm{u}_{1}^{\top}\bm{u}_{2}=0$ [103]. $0=\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{2}-\lambda_{2}\bm{u}_{1}^{\top}\bm{u}_{2}+\eta\bm{u}_{1}^{\top}\bm{u}_{1}=\eta\bm{u}_{1}^{\top}\bm{u}_{1}=\eta$ resulting in $\eta=0$. Thus, we obtain $\bm{\varSigma}\bm{u}_{2}=\lambda_{2}\bm{u}_{2}$ and $\bm{u}_{2}$ is also the eigenvector of the covariance matrix $\bm{\varSigma}$, explaining variance $\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{2}=\lambda_{2}.$ The same procedure is applied to obtain further eigenvectors (PCs).
@@ B.2 Minimum reconstruction error formulation
$=$ $\sum_{\bm{x}}p\left(\bm{x}\right)\left|\sum_{i=1}^{n}\left[r_{i}\left(\bm{x}\right)-f_{i}\left(\bm{x}\right)\right]\bm{u}_{i}\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\left.+\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-g_{i}\right]\bm{u}_{i}\right|^{2}$
@@
$+\sum_{\bm{x}}p\left(\bm{x}\right)\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-g_{i}\right]^{2},$ (54)
@@
$\mathcal{L}\coloneqq\underbrace{\sum_{i=n+1}^{N}\bm{u}_{i}^{\top}\bm{\varSigma}\bm{u}_{i}}_{L}+\sum_{i=n+1}^{N}\lambda_{i}\left(1-\bm{u}_{i}^{\top}\bm{u}_{i}\right).$ (58)
@@ B.3.1 Classical MDS
$\bm{G}=\bm{W}\bm{\varOmega}\bm{W}^{\top},$ (62)
where $\bm{W}\coloneqq\begin{bmatrix}\bm{w}_{1}&\cdots&\bm{w}_{M}\end{bmatrix}\in\mathbb{R}^{M\times M}$ is the matrix with columns being eigenvectors $\{\bm{w}_{i}\}$ and $\bm{\varOmega}\coloneqq\operatorname{diag}\left(\omega_{1},\ldots,\omega_{M}\right)\in\mathbb{R}_{\geq 0}^{M\times M}$ is the diagonal matrix of eigenvalues $\left\{\omega_{i}\right\}$ in descending order $\omega_{1}\geq\cdots\geq\omega_{M}\geq 0$. Noticing that the Gram matrix is the inner product matrix, we can express the Gram matrix as a multiplication of matrices
@@
where $\bm{X}\coloneqq\begin{bmatrix}\bm{x}_{1}&\cdots&\bm{x}_{M}\end{bmatrix}\in\mathbb{R}^{N\times M}$ is the data matrix, with each column being a data point. Then, the rank of the Gram matrix becomes $\operatorname{rank}\left(\bm{G}\right)=\operatorname{rank}\left(\bm{X}^{\top}\bm{X}\right)=\operatorname{rank}\left(\bm{X}\right)=N$. Because the Gram matrix $\bm{G}$ is symmetric and positive semi-definite of rank $N$, it has $N$ non-negative eigenvalues and $M-N$ zero eigenvalues: $\omega_{1}\geq\cdots\geq\omega_{N}>\omega_{N+1}=\cdots=\omega_{M}=0$. We then rewrite the eigendecomposition of the Gram matrix [Eq. (62)] as
$\bm{G}=$ $\tilde{\bm{W}}\tilde{\bm{\varOmega}}\tilde{\bm{W}}^{\top}$ (64)
$=$ $\tilde{\bm{W}}\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{W}}^{\top},$ (65)
with reduced eigenvalue diagonal matrix $\tilde{\bm{\varOmega}}\coloneqq\operatorname{diag}\left(\omega_{1},\ldots,\omega_{N}\right)\in\mathbb{R}_{\geq 0}^{N\times N}$ and reduced eigenvector matrix $\tilde{\bm{W}}\coloneqq\begin{bmatrix}\bm{w}_{1}&\cdots&\bm{w}_{N}\end{bmatrix}\in\mathbb{R}^{M\times N}$ which exclude the zero eigenvalues and corresponding eigenvectors, respectively.
@@
$\bm{X}\bm{C}=\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{W}}^{\top}.$ (67)
To approximate the original data points by the lower-dimensional coordinates while maximally preserving the pairwise distances, one can use the desired number of the first rows of $\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{W}}^{\top}$. For example, if one wants to approximate the original $N$-dimensional data point by two-dimensional coordinates, classical MDS provides a set of scaled coordinates $\bm{R}=\begin{bmatrix}\bm{r}_{1}&\cdots&\bm{r}_{M}\end{bmatrix}=\begin{bmatrix}\sqrt{\omega_{1}}\bm{w}_{1}&\sqrt{\omega_{2}}\bm{w}_{2}\end{bmatrix}^{\top}\in\mathbb{R}^{2\times M},$ where $\bm{r}_{i}$ is the two-dimensional scaled coordinate of the $i$th data point. Replacing original coordinates with principal coordinates is also called principal coordinate analysis (PCO or PCoA) [121].
@@ B.3.2 Classical MDS is equivalent to PCA
$\underbrace{\frac{1}{M}\bm{X}\bm{C}\bm{X}^{\top}}_{\bm{\varSigma}}\left(\bm{X}\bm{C}\bm{w}_{i}\right)=\frac{\omega_{i}}{M}\left(\bm{X}\bm{C}\bm{w}_{i}\right).$ (69)
@@
$\bm{u}_{i}=\frac{1}{\sqrt{\omega_{i}}}\bm{X}\bm{C}\bm{w}_{i}.$ (72)
Following Eq. (68), or $\bm{w}_{i}^{\top}\bm{C}\bm{X}^{\top}\bm{X}\bm{C}\bm{w}_{i}=\omega_{i}$, the normalization factor $\frac{1}{\sqrt{\omega_{i}}}$ is introduced. The PC score by PC loading $\bm{u}_{i}$ and the $i$th element of principal coordinates $\sqrt{\omega_{i}}\bm{w}_{i}^{\top}$ are equivalent:
$\bm{u}_{i}^{\top}\bm{X}\bm{C}=$ $\frac{1}{\sqrt{\omega_{i}}}\bm{w}_{i}^{\top}\bm{C}\bm{X}^{\top}\bm{X}\bm{C}$
$=$ $\sqrt{\omega_{i}}\bm{w}_{i}^{\top}.$ (73)
@@ B.3.3 Weighted classical MDS as PCA
$\underbrace{\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}}_{\bm{G}}\bm{w}_{i}=\omega_{i}\bm{w}_{i}.$
@@
$\underbrace{\bm{X}\acute{\bm{C}}\bm{P}\acute{\bm{C}}^{\top}\bm{X}^{\top}}_{\bm{\varSigma}}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{w}_{i}\right)=\omega_{i}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{w}_{i}\right).$
@@
$\bm{u}_{i}=\frac{1}{\sqrt{\omega_{i}}}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{w}_{i},$
@@
$\bm{u}_{i}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}=$ $\frac{1}{\sqrt{\omega_{i}}}\bm{w}_{i}^{\top}\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}$
$=$ $\sqrt{\omega_{i}}\bm{w}_{i}^{\top}.$

### v4

@@ VIII Conclusions
Our contributions are summarized as follows: (1) We introduce the Hamming and fractal projection of hypercubes in Sec. II. (2) We propose interpreting the biplot vectors of PCA as the vectors of the orthogonal projection of hypercubes in Sec. III. (3) We find that the PC1 loading of PCA on hypercubes is equivalent to the weighted hypercubic vertices in Sec. IV, and validate it analytically and numerically in Sec. V. (4) We reveal numerically and analytically that the quality of the orthogonal projection of hypercubes tends to be worse around the origin of the projected space in Sec. VI. (5) We visualize the hypercubic energy landscapes of finite artificial spin-ice systems and extract physical interpretations—particularly focusing on correlated spin flips—in Sec. VII.
Although we have demonstrated that PCA yields reproducible, interpretable, and automatic projections of hypercubes, there are two remaining challenges: (1) analytically predicting the eigenvectors of the covariance matrix for non-bipolar probability distributions, and (2) the overlap of hypercubic vertices and hypercubic edges in the projected space. The first challenge is not yet resolved in this study, but we believe that the symmetric outer product matrix [the second term of Eq. (36)] is key to addressing this problem. We present some nontrivial examples in Fig. 4(a) and Supplemental Material [16], where the PC loading is not proportional to the most heavily weighted hypercubic vertex. Perturbation theory (Appendix C) is promising to address this challenge. Turning to the second challenge, many hypercubic vertices and edges tend to overlap in the projected space, especially near the origin. As we show numerically and analytically in Sec. VI, PCA locates numerous hypercubic vertices around the origin of the projected space, and those overlapped hypercubic vertices can increase the error of the projection. We, however, analytically know these limitations of PCA beforehand: we can interpret the resulting plot while grasping the limitations, unlike other nonlinear dimensionality reduction methods, where such analytical properties are not obvious and might be more difficult to understand. When constructing the hypercubic energy landscape and probability flux in Sec. VII, we mitigate the overlap problem by projecting only a limited subset of hypercubic vertices and edges. A similar strategy may be employed to bypass this issue in other applications.
H. S. M. Coxeter mentioned that illustrations of high-dimensional objects have psychological and artistic merit [1]. In this study, we initiate unveiling the new merit—or possibility—of projecting high-dimensional hypercubes: interpreting binary data through data-driven visualization. As Anscombe’s quartet [124] shows hidden patterns not in statistics but in graphs [125], informative orthogonal projection using PCA might lead to fresh understanding, interpretation, and discovery [126, 127, 128] in high-dimensional binary data across sciences.
Calculations and visualizations of this work were performed using open-source python [129] libraries: matplotlib [130], networkx [131], numpy [132], and scipy [133]. The color map of some figures are generated by colorcet [134]. All data and code are available online from Zenodo [135].
@@ Acknowledgements.
Special thanks go to Takashi Horiike, Kennichi Hagiwara, Nozomu Imai, Takumi Nishina, and Hiroto Murata for fruitful discussions at the early stage of this research. Y.H. thank Mizuho Aoki, Johanna von Drachenfels, Anton G. Golles, and Henri Schmidt for their valuable comments on this manuscript. This work was supported by KAKENHI Grant Number 22H00406 and 24H00061 of the Japan Society for the Promotion of Science. Y.H. is supported by JST SPRING, Grant Number JPMJSP2125 “THERS Make New Standards Program for the Next Generation Researchers”.
Y.H. performed conceptualization, data curation, formal analysis, investigation, software development, validation, and visualization; Y.H. partially contributed to the funding acquisition; Y.H. and S.F. developed the methodology; S.F. provided computational resources; Y.H. wrote the original draft; and Y.H. and S.F. reviewed and edited the manuscript.
@@ Appendix A Constructing hypercubes
A hypercube is constructed as follows [1, 136, 4]. Suppose one has a zero-dimensional point, one obtains a one-dimensional line by moving the point in a direction. By moving the line in a direction but not along its line, one can obtain a two-dimensional parallelogram. Again, moving the parallelogram in a direction but not along its plane, one creates a three-dimensional parallelepiped. By repeating this translation process $N$ times, one obtains an $N$-dimensional parallelepiped or a parallelotope with $2^{N}$ vertices and $\frac{2^{N}N}{2}$ edges.141414 Each of the $2^{N}$ vertices has $N$ edges. To avoid double counting, one needs to divide them by two. The unit vectors of each dimension correspond to the direction of the translation, i.e., edges created by the translation are parallel to the directions of unit vectors.
@@ B.1 Maximum projection variance formulation
Multiplying $\bm{u}_{1}^{\top}$ to both sides of the equation from the left side yields151515 $\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{2}=\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{1}=\bm{u}_{2}^{\top}\lambda_{1}\bm{u}_{1}=\lambda_{1}\bm{u}_{2}^{\top}\bm{u}_{1}=0$ because of the orthogonality between $\bm{u}_{1}$ and $\bm{u}_{2}$, $\bm{u}_{1}^{\top}\bm{u}_{2}=0$ [93]. $0=\bm{u}_{1}^{\top}\bm{\varSigma}\bm{u}_{2}-\lambda_{2}\bm{u}_{1}^{\top}\bm{u}_{2}+\eta\bm{u}_{1}^{\top}\bm{u}_{1}=\eta\bm{u}_{1}^{\top}\bm{u}_{1}=\eta$ resulting in $\eta=0$. Thus, we obtain $\bm{\varSigma}\bm{u}_{2}=\lambda_{2}\bm{u}_{2}$ and $\bm{u}_{2}$ is also the eigenvector of the covariance matrix $\bm{\varSigma}$, explaining variance $\bm{u}_{2}^{\top}\bm{\varSigma}\bm{u}_{2}=\lambda_{2}.$ The same procedure is applied to obtain further eigenvectors (PCs).
@@ B.2 Minimum reconstruction error formulation
$=\sum_{\bm{x}}p\left(\bm{x}\right)\left|\sum_{i=1}^{n}\left[r_{i}\left(\bm{x}\right)-f_{i}\left(\bm{x}\right)\right]\bm{u}_{i}+\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-g_{i}\right]\bm{u}_{i}\right|^{2}$
@@
$\quad+\sum_{\bm{x}}p\left(\bm{x}\right)\sum_{i=n+1}^{N}\left[r_{i}\left(\bm{x}\right)-g_{i}\right]^{2},$ (98)
@@
$\mathcal{L}\coloneqq\underbrace{\sum_{i=n+1}^{N}\bm{u}_{i}^{\top}\bm{\varSigma}\bm{u}_{i}}_{=L}+\sum_{i=n+1}^{N}\lambda_{i}\left(1-\bm{u}_{i}^{\top}\bm{u}_{i}\right).$ (102)
@@ B.3.1 Classical MDS
$\bm{G}=\bm{Y}\bm{\varOmega}\bm{Y}^{\top},$ (106)
where $\bm{Y}\coloneqq\begin{bmatrix}\bm{y}_{1}&\cdots&\bm{y}_{M}\end{bmatrix}\in\mathbb{R}^{M\times M}$ is the matrix with columns being eigenvectors $\{\bm{y}_{i}\}$ and $\bm{\varOmega}\coloneqq\operatorname{diag}\left(\omega_{1},\ldots,\omega_{M}\right)\in\mathbb{R}_{\geq 0}^{M\times M}$ is the diagonal matrix of eigenvalues $\left\{\omega_{i}\right\}$ in descending order $\omega_{1}\geq\cdots\geq\omega_{M}\geq 0$. Noticing that the Gram matrix is the inner product matrix, we can express the Gram matrix as a multiplication of matrices
@@
where $\bm{X}\coloneqq\begin{bmatrix}\bm{x}_{1}&\cdots&\bm{x}_{M}\end{bmatrix}\in\mathbb{R}^{N\times M}$ is the data matrix, with each column being a data point.161616 Note that our definition of the data matrix is the transpose of the convention commonly used in the statistics literature. Then, the rank of the Gram matrix becomes $\operatorname{rank}\left(\bm{G}\right)=\operatorname{rank}\left(\bm{X}^{\top}\bm{X}\right)=\operatorname{rank}\left(\bm{X}\right)=N$. Because the Gram matrix $\bm{G}$ is symmetric and positive semi-definite of rank $N$, it has $N$ non-negative eigenvalues and $M-N$ zero eigenvalues: $\omega_{1}\geq\cdots\geq\omega_{N}>\omega_{N+1}=\cdots=\omega_{M}=0$. We then rewrite the eigendecomposition of the Gram matrix [Eq. (106)] as
$\bm{G}$ $=\tilde{\bm{Y}}\tilde{\bm{\varOmega}}\tilde{\bm{Y}}^{\top}$ (108)
$=\tilde{\bm{Y}}\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{Y}}^{\top},$ (109)
with reduced eigenvalue diagonal matrix $\tilde{\bm{\varOmega}}\coloneqq\operatorname{diag}\left(\omega_{1},\ldots,\omega_{N}\right)\in\mathbb{R}_{\geq 0}^{N\times N}$ and reduced eigenvector matrix $\tilde{\bm{Y}}\coloneqq\begin{bmatrix}\bm{y}_{1}&\cdots&\bm{y}_{N}\end{bmatrix}\in\mathbb{R}^{M\times N}$ which exclude the zero eigenvalues and corresponding eigenvectors, respectively.
@@
$\bm{X}\bm{C}=\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{Y}}^{\top}.$ (111)
To approximate the original data points by the lower-dimensional coordinates while maximally preserving the pairwise distances, one can use the desired number of the first rows of $\tilde{\bm{\varOmega}}^{\frac{1}{2}}\tilde{\bm{Y}}^{\top}$. For example, if one wants to approximate the original $N$-dimensional data point by two-dimensional coordinates, classical MDS provides a set of scaled coordinates $\bm{R}=\begin{bmatrix}\bm{r}_{1}&\cdots&\bm{r}_{M}\end{bmatrix}=\begin{bmatrix}\sqrt{\omega_{1}}\bm{y}_{1}&\sqrt{\omega_{2}}\bm{y}_{2}\end{bmatrix}^{\top}\in\mathbb{R}^{2\times M},$ where $\bm{r}_{i}$ is the two-dimensional scaled coordinate of the $i$th data point. Replacing original coordinates with principal coordinates is also called principal coordinate analysis (PCO or PCoA) [110].
@@ B.3.2 Classical MDS is equivalent to PCA
$\underbrace{\frac{1}{M}\bm{X}\bm{C}\bm{X}^{\top}}_{=\bm{\varSigma}}\left(\bm{X}\bm{C}\bm{y}_{i}\right)=\frac{\omega_{i}}{M}\left(\bm{X}\bm{C}\bm{y}_{i}\right).$ (113)
@@
$\bm{u}_{i}=\frac{1}{\sqrt{\omega_{i}}}\bm{X}\bm{C}\bm{y}_{i}.$ (116)
Following Eq. (112), or $\bm{y}_{i}^{\top}\bm{C}\bm{X}^{\top}\bm{X}\bm{C}\bm{y}_{i}=\omega_{i}$, the normalization factor $\frac{1}{\sqrt{\omega_{i}}}$ is introduced. The PC score by PC loading $\bm{u}_{i}$ and the $i$th element of principal coordinates $\sqrt{\omega_{i}}\bm{y}_{i}^{\top}$ are equivalent:
$\bm{u}_{i}^{\top}\bm{X}\bm{C}$ $=\frac{1}{\sqrt{\omega_{i}}}\bm{y}_{i}^{\top}\bm{C}\bm{X}^{\top}\bm{X}\bm{C}$
$=\sqrt{\omega_{i}}\bm{y}_{i}^{\top}.$ (117)
@@ B.3.3 Weighted classical MDS as PCA
$\underbrace{\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}}_{=\bm{G}}\bm{y}_{i}=\omega_{i}\bm{y}_{i}.$
@@
$\underbrace{\bm{X}\acute{\bm{C}}\bm{P}\acute{\bm{C}}^{\top}\bm{X}^{\top}}_{=\bm{\varSigma}}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{y}_{i}\right)=\omega_{i}\left(\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{y}_{i}\right).$
@@
$\bm{u}_{i}=\frac{1}{\sqrt{\omega_{i}}}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}\bm{y}_{i},$
@@
$\bm{u}_{i}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}$ $=\frac{1}{\sqrt{\omega_{i}}}\bm{y}_{i}^{\top}\bm{P}^{\frac{1}{2}}\acute{\bm{C}}^{\top}\bm{X}^{\top}\bm{X}\acute{\bm{C}}\bm{P}^{\frac{1}{2}}$
$=\sqrt{\omega_{i}}\bm{y}_{i}^{\top}.$